OpenStudy (anonymous):

Can anyone calculate the speed and force of an object of a one ton flat object travelling on a 15 degrees slope over 5 minutes? Assume there is no friction.

OpenStudy (anonymous):

Not without more data, no.

OpenStudy (anonymous):

fist find accelaration due to gravity at 15 degrees it is 9.46 then use the distance formula d=.5*a*t^2 where t is in seconds so d=.5*9.46*(60*5)^2= 4259.73 meter.

OpenStudy (anonymous):

now we use kenergy=.5m*v^2 and and potential energy = f*d=(a*m)*d So pe= 4259.73*(9.46*1000) = 40297045.80 joules so now we plug that into kenergy equations to get speed 40297045.80 = (.5*1000)*v^2 solve for v to get velocity 283.89 m/s so v is 283.89 m/s F is 9460.00 Newtons

OpenStudy (anonymous):

If I am wrong then use acceleration as 2.53 Sounds more reasonalbe. and do same steps with a as 2.53 instea of 9.46

OpenStudy (anonymous):

with a - 2.53 f = 2435.95 N and V is 760 m/s

OpenStudy (anonymous):

*a =2.53

OpenStudy (anonymous):

We also can use dynamical and then kinematical approach, for solving of this problem.  Let be F the component of gravitational force mg along the direction of motion of the object, no other forces are involved except normal force of reaction in which we are not interested in, since there is no friction. We get F = mg sin(15 dgr) This is the force you've been asked for. Now using the second law of Newton we find a = F/m and then using equation of kinematics for velocity for motion with constant acceleration, we get speed v = v0+at v0=0 because, by default we assume that at the beginning the object was at rest. So, v = at and that's it.